Circular No. 6778

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE SUPERNOVA IN UGC 4107 S. Nakano, Sumoto, Japan, reports the discovery by Yasuo Sano, Nay-

oro, Hokkaido, of a possible supernova (R = 16.7) on ten CCD frames taken ∼ Nov. 25.582 UT with a 0.28-m f/6.3 reflector (+ R filter). The object is located at α = 7^{h}57^{m}02^{s}.82, δ = +49^{o}33′40′′.2 (equinox 2000.0), which is ∼ 10′′ east and 20′′ south of the center of UGC 4107. R. Kushida, Yatsugatake South Base Observatory, provides the following position end figures (measured by Y. Kushida) from an unfiltered CCD image that she obtained on Nov. 26.81 (when the supernova suspect was at mag 16.1) with a 0.40-m f/5 reflector: 02^{s}.87, 41′′.3. There is no star at this position on the Digital Sky Survey.

P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics, communicate: “P. Berlind obtained a spectrogram (range 370–750 nm) of Sano’s supernova candidate in UGC 4107 on Nov. 26.4 UT with the Whipple Observatory’s 1.5-m Tillinghast telescope. The spectrum is unusual, showing a continuum with broad (50- to 100-nm) features. The continuum peaks near 510 nm and shows local maxima at 440 and 620 nm. Narrow emission lines from UGC 4107 are also seen and give a redshift of 3400 km/s. The spectrum is clearly not that of a normal star, but from the present evidence, it is not yet possible to conclude that the object is a supernova. The spectrum is available at http://cfa-

www.harvard.edu/cfa/oir/Research/supernova/RecentSN.html.”

NOVA SCORPII 1997 M. Orio and E. Trussoni, Turin Observatory; S. Balman, Middle East-

ern Technical University, Ankara; H. Ogelman and J. Gallagher, University of Wisconsin; D. de Martino, Naples Observatory; M. Della Valle, Padua University; C. Gonzalez-Riestra, European Space Agency; and P. Selvelli, National Research Council of Italy, Trieste, report: “N Sco 1997 (cf. IAUC 6678, 6679) was detected as an x-ray source by the SAX LECS and MECS instruments with countrates 0.0658 ± 0.0021 and 0.0190 ± 0.0007 counts/s, respectively, on Sept. 23–24 UT in the two instruments. The unabsorbed flux from the source is 4.5 × 10^{−12} erg cm^{−2} s^{−1} in the band 0.1–10 keV, assuming that the emitting material is a two-component thermal plasma with a column of neutral hydrogen, N_{H}, ∼ 2 × 10^{21} cm^{−2}; this assumption is consistent with a shocked-shell model.”

1997 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green